# Data-Driven Reduced-Order Models for Port-Hamiltonian Systems with Operator Inference

### v1

@@ 5.2 Linear Case: Mass-Spring-Damper System
Next, we test the influence of the regularization parameter $\lambda_{R}$ on the optimization problem and the resulting effects on the pH-OpInf-R ROM. When $\lambda_{R}=0$, we found that ${\mathbf{B}}_{r}$ solved from equation (14) using least squares is poorly conditioned, leading to unstable ROM simulations. Taking $r=5,10,15$ and $20$ and selecting eight values of $\lambda_{R}$ from the set $\{10^{-14},10^{-13},\dots,10^{-7}\}$, we solve (14) and (15), respectively. The optimization errors $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ are shown in Figure 3. For $r=5$ and $10$, the optimization errors are relatively steady as $\lambda_{R}$ varies. However, for larger values of $r=15$ and $20$, as $\lambda_{R}$ increases, both $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ initially decrease (with $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ remaining unchanged for $r=15$), but after $\lambda_{R}=10^{-11}$, both begin to increase. Therefore, in subsequent experiments, we select $\lambda_{R}=10^{-11}$ in pH-OpInf-R.
@@
Setting $T_{\text{FOM}}=T_{\text{ROM}}=10$ and gradually increasing the ROM dimension $r$ by $5$, from $5$ to $50$, we simulate the pH-OpInf-ROMs. The associated numerical errors defined in (30)–(35) are plotted in Figure 4: the left column corresponds to pH-OpInf-W, and the right column corresponds to pH-OpInf-R. From the top row, we observe that in both cases, the projection error $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ continues to decay as $r$ increases, eventually reaching the machine precision beyond $r=25$. Meanwhile, the optimization error $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ decreases to around $10^{-5}$ as $r$ increases to $20$ and remains nearly steady. Consequently, the state approximation error $\mathcal{E}_{\mathbf{x}}$ is dominated by $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and exhibits similar behavior. From the middle row, we find that the optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ from pH-OpInf-R is smaller than that from pH-OpInf-W when $r>5$, but in both cases, $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ is much smaller than the state approximation error $\mathcal{E}_{\mathbf{x}}$. Thus, the output approximation error $\mathcal{E}_{\mathbf{y}}$ is dominated by $\mathcal{E}_{\mathbf{x}}$ and displays similar behavior. In the bottom row, $\mathcal{E}_{\mathbf{y}}$ from the pH-OpInf-ROMs is compared with that of the SP-G-ROM. The pH-OpInf-ROM, based on either pH-OpInf-W or pH-OpInf-R, attains better results when $r<25$. Beyond that, SP-G-ROM achieves better approximations. Therefore, in case when the computation budget is a concern, a ROM with a lower dimension is preferable, and the pH-OpInf-ROM is the better choice. Furthermore, the time evolution of the output and approximate Hamiltonian from the pH-OpInf-ROMs is shown in Figure 5. We observe that the reduced-order approximations of both the output and the Hamiltonian, using either pH-OpInf-W or pH-OpInf-R, closely match those from the FOM simulation when $r\geq 10$.
@@ 5.3 Nonlinear Case: Toda Lattice Model
For this nonlinear pH system, we first consider pH-OpInf-ROM without using the hyper-reduction. To illustrate the error estimation given in Corollary 2.1, we choose a small time step size $\Delta t=0.0025$ in both FOM and ROM simulations, set $T_{\text{FOM}}=T_{\text{ROM}}=50$ and vary the dimension $r$ of the pH-OpInf-ROM from 10 to 100, increasing incrementally by 10. The associated numerical errors defined in (30)–(35) are plotted in Figure 10: the left column corresponds to pH-OpInf-W, and the right column corresponds to pH-OpInf-R. The top row displays $\mathcal{E}_{\mathbf{x}}$, $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$. For both pH-OpInf-W and pH-OpInf-R, $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ decreases as the dimension $r$ increases up to $r=60$, beyond which it saturates. However, $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ continues to decline until $r=90$, for which it reaches machine precision. As a result, the state approximation error $\mathcal{E}_{\mathbf{x}}$ is dominated by $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and follows a similar trend. The output approximation error $\mathcal{E}_{\mathbf{y}}$ is shown alongside $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ and $\mathcal{E}_{\mathbf{x}}$ in the middle row, from which we observe that the optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ from pH-OpInf-R is smaller than that from pH-OpInf-W when $r$ is bigger than $40$. In both cases, $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ is smaller than $\mathcal{E}_{\mathbf{x}}$, thus the output approximation error $\mathcal{E}_{\mathbf{y}}$ is primarily influenced by $\mathcal{E}_{\mathbf{x}}$ and follows a similar pattern. The bottom row in Figure 10 compares the approximation errors of pH-OpInf-ROM with the intrusive SP-G-ROM. The results show that the pH-OpInf-ROM, based on either pH-OpInf-W or pH-OpInf-R, achieves better approximations than the instrusive SP-G-ROM for dimensions up to a fairly large dimension $r=80$.
@@
The time evolution of the output and approximate Hamiltonian from the pH-OpInf-ROMs is presented in Figure 11. We observe that the reduced-order approximations, using either pH-OpInf-W or pH-OpInf-R, get more accurate as $r$ increases, and when $r=50$, they closely align with those from the FOM simulation. Similar to the linear case in Section 5.2, although both achieve good approximations, the pH-OpInf-R attains a smaller optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$. Therefore, we use pH-OpInf-R in the subsequent experiments.
@@
and ${\mathbf{c}}=[1,1,\dots,1]^{\intercal}\in\mathbb{R}^{N_{0}}$ with $h_{1}=\exp(q_{1}-q_{2})-q_{1}-N_{0}$, and $h_{i}=\exp(q_{i}-q_{i+1})$ for $i=2,3,\dots,N_{0}-1$ and $h_{N_{0}}=\exp(q_{N_{0}})$. Choosing $r=20$ and $50$, respectively, we simulate the pH-OpInf-DEIM using $m$ interpolation points in DEIM and compare it with the pH-OpInf-ROM of the same dimensions. The numerical errors in state and output approximations are listed in Table 3, together with the DEIM interpolation error $\mathcal{E}_{\text{DEIM}}$. Overall, $\mathcal{E}_{\text{DEIM}}$ decreases quickly as $m$ increases. For both $r=20$ and $50$, it reaches machine precision when $m=60$, and the pH-OpInf-DEIM results coincide with those of pH-OpInf-ROM. This observation matches our theoretical results, as stated in Theorems 1 and 2, because, comparing with pH-OpInf-ROM, the error bound of pH-OpInf-DEIM involves the additional term $\mathcal{E}_{\text{DEIM}}$. When $\mathcal{E}_{\text{DEIM}}$ is sufficiently small, errors from both ROMs should become identical.
Table 3: Toda Lattice Model: Errors of pH-OpInf-ROM with dimension $r$ and pH-OpInf-DEIM with the same dimension and $m$ DEIM interpolation points.
@@
$\mathcal{E}_{\mathbf{x}}$ 6.042 ×$\times 10^{-1}$ 6.355×$\times 10^{-1}$ 5.825×$\times 10^{-1}$ 6.042×$\times 10^{-1}$ 6.042×$\times 10^{-1}$
$\mathcal{E}_{\mathbf{y}}$ 4.032×$\times 10^{-2}$ 4.794×$\times 10^{-2}$ 4.028×$\times 10^{-2}$ 4.032×$\times 10^{-2}$ 4.032×$\times 10^{-2}$
$\mathcal{E}_{\text{DEIM}}$ 6.104×$\times 10^{-2}$ 8.038×$\times 10^{-4}$ 1.037×$\times 10^{-7}$ 2.798×$\times 10^{-14}$
pH-OpInf-ROM with $r=50$ pH-OpInf-DEIM with $r=50$
$m=45$ $m=50$ $m=55$ $m=60$
$\mathcal{E}_{\mathbf{x}}$ 2.354×$\times 10^{-4}$ 8.339×$\times 10^{-3}$ 2.301×$\times 10^{-4}$ 2.353×$\times 10^{-4}$ 2.354×$\times 10^{-4}$
$\mathcal{E}_{\mathbf{y}}$ 7.799×$\times 10^{-6}$ 1.154×$\times 10^{-3}$ 7.733×$\times 10^{-6}$ 7.799×$\times 10^{-6}$ 7.799×$\times 10^{-6}$
$\mathcal{E}_{\text{DEIM}}$ 3.436×$\times 10^{-3}$ 3.415×$\times 10^{-6}$ 4.756×$\times 10^{-9}$ 1.994×$\times 10^{-13}$
@@ 6 Conclusions
The numerical results show that both pH-OpInf-W and pH-OpInf-R with carefully chosen optimization parameters find reduced-order operators that preserve the appropriate structures. Generally, the overall performance of these two approaches is comparable, since the pH-OpInf-ROM constructed using either approach yields accurate outputs in response to inputs. However, pH-OpInf-R achieves a smaller optimization error for the output, $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$, than pH-OpInf-W, which makes it the preferred choice. Furthermore, for nonlinear pH systems, the pH-OpInf-DEIM model using sufficient DEIM interpolation points is able to achieve the same performance at a reduced computational cost.
@@ Acknowledgement
Y.G. was partially supported by a SPARC grant from the University of South Carolina. L.J. was partially supported by the U.S. Department of Energy under award numbers DE-SC0022254 and DE-SC0025527. B.K. and Z.W. were supported by U.S. Office of Naval Research under award number N00014-22-1-2624. Z.W. was partially supported by U.S. National Science Foundation under award numbers DMS-2012469, DMS-2038080, DMS-2245097 and an ASPIRE grant from the Office of the Vice President for Research at the University of South Carolina.

### v2

@@ 5.2 Linear Case: Mass-Spring-Damper System
Next, we test the influence of the regularization parameter $\lambda_{R}$ on the optimization problem and the resulting effects on the pH-OpInf-R ROM. When $\lambda_{R}=0$, we find that the solution ${\mathbf{B}}_{r}$ from equation (14) using least squares is poorly conditioned, leading to unstable ROM simulations. Taking $r=5,10,15$ and $20$ and selecting eight values of $\lambda_{R}$ from the set $\{10^{-14},10^{-13},\dots,10^{-7}\}$, we solve (14) and (15), respectively. The optimization errors $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ are shown in Figure 3. For $r=5$ and $10$, the optimization errors are relatively steady as $\lambda_{R}$ varies. However, for larger values of $r=15$ and $20$, as $\lambda_{R}$ increases, both $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ initially decrease (with $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ remaining unchanged for $r=15$), but after $\lambda_{R}=10^{-11}$, both begin to increase. Therefore, in subsequent experiments, we select $\lambda_{R}=10^{-11}$ in pH-OpInf-R.
@@
Setting $T_{\text{FOM}}=T_{\text{ROM}}=10$ and gradually increasing the ROM dimension $r$ by $5$, from $5$ to $50$, we simulate the pH-OpInf-ROMs. In this case, the singular values of the snapshot matrix ${\mathbf{X}}$ decay quickly: the first 5 POD modes capture 95.24% of the snapshot energy, while the first 10 modes capture 99.99%. The associated numerical errors defined in (31)–(36) are plotted in Figure 4: the left column corresponds to pH-OpInf-W, and the right column corresponds to pH-OpInf-R. Note that the magnitude of the state vectors is 0.661, while that of the output is 0.083. From the top row, we observe that in both cases, the projection error $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ continues to decay as $r$ increases, eventually reaching the machine precision beyond $r=25$. Meanwhile, the optimization error $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ decreases to around $10^{-5}$ as $r$ increases to $20$ and remains nearly steady. Consequently, the state approximation error $\mathcal{E}_{\mathbf{x}}$ is dominated by $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and exhibits similar behavior. From the middle row, we find that the optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ from pH-OpInf-R is smaller than that from pH-OpInf-W when $r>5$, but in both cases, $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ is much smaller than the state approximation error $\mathcal{E}_{\mathbf{x}}$. Thus, the output approximation error $\mathcal{E}_{\mathbf{y}}$ is dominated by $\mathcal{E}_{\mathbf{x}}$ and displays similar behavior. In the bottom row, $\mathcal{E}_{\mathbf{y}}$ from the pH-OpInf-ROMs is compared with that of the SP-G-ROM. The pH-OpInf-ROM, based on either pH-OpInf-W or pH-OpInf-R, attains better results when $r<25$. Beyond that, SP-G-ROM achieves better approximations. Therefore, in case when the computation budget is a concern, a ROM with a lower dimension is preferable, and the pH-OpInf-ROM is the better choice. Furthermore, the time evolution of the output and approximate Hamiltonian from the pH-OpInf-ROMs is shown in Figure 5. We observe that the reduced-order approximations of both the output and the Hamiltonian, using either pH-OpInf-W or pH-OpInf-R, closely match those from the FOM simulation when $r\geq 10$.
@@ 5.3 Nonlinear Case: Toda Lattice Model
For this nonlinear pH system, we first consider pH-OpInf-ROM without using the hyper-reduction. To illustrate the error estimation given in Corollary 2.1, we choose a small time step size $\Delta t=0.0025$ in both FOM and ROM simulations, set $T_{\text{FOM}}=T_{\text{ROM}}=50$ and vary the dimension $r$ of the pH-OpInf-ROM from 10 to 100, increasing incrementally by 10. In this case, the singular values of the snapshot matrix ${\mathbf{X}}$ decay more slowly than those in the mass-spring-damper case: the first 10 POD modes capture 97.97% of snapshot energy, the first 20 capture 99.8%, and the first 30 already capture 99.99%. The associated numerical errors defined in (31)–(36) are plotted in Figure 10: the left column corresponds to pH-OpInf-W, and the right column corresponds to pH-OpInf-R. Note that the magnitude of the state vectors is 2.225, while that of the output is 0.483. The top row displays $\mathcal{E}_{\mathbf{x}}$, $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ and $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$. For both pH-OpInf-W and pH-OpInf-R, $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ decreases as the dimension $r$ increases up to $r=60$, beyond which it saturates. However, $\mathcal{E}_{\text{proj}_{\mathbf{x}}}$ continues to decline until $r=90$, for which it reaches machine precision. As a result, the state approximation error $\mathcal{E}_{\mathbf{x}}$ is dominated by $\mathcal{E}_{\text{opt}_{\mathbf{x}}}$ and follows a similar trend. The output approximation error $\mathcal{E}_{\mathbf{y}}$ is shown alongside $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$ and $\mathcal{E}_{\mathbf{x}}$ in the middle row, from which we observe that the optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ from pH-OpInf-R is smaller than that from pH-OpInf-W when $r$ is bigger than $40$. In both cases, $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$ is smaller than $\mathcal{E}_{\mathbf{x}}$, thus the output approximation error $\mathcal{E}_{\mathbf{y}}$ is primarily influenced by $\mathcal{E}_{\mathbf{x}}$ and follows a similar pattern. The bottom row in Figure 10 compares the approximation errors of pH-OpInf-ROM with the intrusive SP-G-ROM. The results show that the pH-OpInf-ROM, based on either pH-OpInf-W or pH-OpInf-R, achieves better approximations than the instrusive SP-G-ROM for dimensions up to a fairly large dimension $r=80$.
@@
The time evolution of the output and approximate Hamiltonian from the pH-OpInf-ROMs is presented in Figure 11 when $r=20$, $40$ and $60$, respectively. We observe that the reduced-order approximations, using either pH-OpInf-W or pH-OpInf-R, get more accurate as $r$ increases, and when $r=40$, they closely align with those from the FOM simulation. Similar to the linear case in Section 5.2, although both achieve good approximations, the pH-OpInf-R attains a smaller optimization error $\mathcal{E}_{\text{opt}_{{\mathbf{y}}}}$. Therefore, we use pH-OpInf-R in the subsequent experiments.
@@
and ${\mathbf{c}}=[1,1,\dots,1]^{\intercal}\in\mathbb{R}^{N_{0}}$ with $h_{1}=\exp(q_{1}-q_{2})-q_{1}-N_{0}$, and $h_{i}=\exp(q_{i}-q_{i+1})$ for $i=2,3,\dots,N_{0}-1$ and $h_{N_{0}}=\exp(q_{N_{0}})$. Choosing $r=20$ and $60$, respectively, we simulate the pH-OpInf-DEIM using $m$ interpolation points in DEIM and compare it with the pH-OpInf-ROM of the same dimensions. The numerical errors in state and output approximations are listed in Table 3, together with the DEIM interpolation error $\mathcal{E}_{\text{DEIM}}$. Overall, $\mathcal{E}_{\text{DEIM}}$ decreases quickly as $m$ increases. For both $r=20$ and $60$, its precision reaches $10^{-11}$ or higher when $m=60$, and the pH-OpInf-DEIM results coincide with those of pH-OpInf-ROM. This observation matches our theoretical results, as stated in Theorems 1 and 2, because, comparing with pH-OpInf-ROM, the error bound of pH-OpInf-DEIM involves the additional term $\mathcal{E}_{\text{DEIM}}$. When $\mathcal{E}_{\text{DEIM}}$ is sufficiently small, errors from both ROMs should become identical.
@@
$\mathcal{E}_{\mathbf{x}}$ 6.042 ×$\times 10^{-1}$ 6.494×$\times 10^{-1}$ 5.958×$\times 10^{-1}$ 6.042×$\times 10^{-1}$ 6.042×$\times 10^{-1}$
$\mathcal{E}_{\mathbf{y}}$ 4.032×$\times 10^{-2}$ 4.400×$\times 10^{-2}$ 4.030×$\times 10^{-2}$ 4.032×$\times 10^{-2}$ 4.032×$\times 10^{-2}$
$\mathcal{E}_{\text{DEIM}}$ 3.985×$\times 10^{-2}$ 7.552×$\times 10^{-4}$ 4.095×$\times 10^{-8}$ 4.585×$\times 10^{-15}$
pH-OpInf-ROM with $r=60$ pH-OpInf-DEIM with $r=60$
$m=50$ $m=55$ $m=60$ $m=65$
$\mathcal{E}_{\mathbf{x}}$ 1.719×$\times 10^{-4}$ 1.745×$\times 10^{-4}$ 1.719×$\times 10^{-4}$ 1.719×$\times 10^{-4}$ 1.719×$\times 10^{-4}$
$\mathcal{E}_{\mathbf{y}}$ 7.748×$\times 10^{-6}$ 7.781×$\times 10^{-6}$ 7.748×$\times 10^{-6}$ 7.748×$\times 10^{-6}$ 7.748×$\times 10^{-6}$
$\mathcal{E}_{\text{DEIM}}$ 1.184×$\times 10^{-5}$ 4.384×$\times 10^{-9}$ 3.478×$\times 10^{-11}$ 1.737×$\times 10^{-13}$
Table 3: Toda Lattice Model: Errors of pH-OpInf-ROM with dimension $r$ and pH-OpInf-DEIM with the same dimension using $m$ DEIM interpolation points.
@@ 6 Conclusions
The numerical results show that both pH-OpInf-W and pH-OpInf-R with carefully chosen optimization parameters find reduced-order operators that preserve the appropriate structures. Generally, the overall performance of these two approaches is comparable, since the pH-OpInf-ROM constructed using either approach yields accurate outputs in response to inputs. However, pH-OpInf-R achieves a smaller optimization error for the output, $\mathcal{E}_{\text{opt}_{\mathbf{y}}}$, than pH-OpInf-W, which makes it the preferred choice. Furthermore, for nonlinear pH systems, the pH-OpInf-DEIM model using sufficient DEIM interpolation points is able to achieve the same performance at a reduced computational cost. We note that alternative hyper-reduction approaches, such as ECSW [52], may also be applied to inferred ROMs when they are nonlinear. We will investigate their performance in future work.
@@ Acknowledgement
Y.G. was partially supported by the SPARC grant from the University of South Carolina under award number CL071-130600-N1400-202-80006259-1. L.J. was partially supported by the U.S. Department of Energy under award numbers DE-SC0022254 and DE-SC0025527. B.K. and Z.W. were supported by U.S. Office of Naval Research under award number N00014-22-1-2624. Z.W. was partially supported by U.S. National Science Foundation under award numbers DMS-2012469, DMS-2038080, DMS-2245097 and an ASPIRE grant from the Office of the Vice President for Research at the University of South Carolina.